OpenStudy (anonymous):

Factor completely 144x2 – 120x + 25

OpenStudy (mertsj):

It's a trinomial square. Do you know what that is?

OpenStudy (anonymous):

no.

OpenStudy (anonymous):

(12x - 5)^2

OpenStudy (hoblos):

a^2 - 2ab +b^2 = (a-b)^2 try to use this

OpenStudy (mertsj):

It is a trinomial that can be factored as a binomial squared.

OpenStudy (hoblos):

oh yeah you are right :D

OpenStudy (mertsj):

For example: x^2 + 6x + 9 = (x+3)^2

OpenStudy (mertsj):

You can get a clue that it might be a trinomial square when you see that the first term which is 144x^2 is a perfect square and the last term which is 25 is a perfect square.

OpenStudy (mertsj):

Yes. You factored it correctly. Good for you!!

OpenStudy (anonymous):

144 and 25 are squares